The two cylindrical rod segments are fixed to the rigid walls such that there is a gap of 0.01in between

them when T1 = 60oF. Determine the larger temperature T2 required in order to just close the gap

without producing stresses in the rods? Each rod has a diameter of 1.25in. Take (alpha)al = 13 x 10-6 in/in oF,

Eal = 104 ksi, (Alpha)cu = 9.4 x 10-6 in/in oF, Ecu = 15 x 103 ksi.




1
Expert's answer
2021-01-06T06:05:40-0500

a)="\Delta Cu + \Delta Al = 0.01"

"L cu D cu (T2-T1) + L al D al (T2-T1) = 0.01"

"T2 = 112.4 F"

b)= "[-(12\times9.4\times10" -6) "- (6\times13\times10" -6)]+"\alpha [ \frac{12} {10^4}+ \frac{6}{1.500^4}]" = 0.01


= 34.87Ksi